In one city, the probability that a person will pass his or her driving test on the first attempt is 0.67. 11 people are selected at random from among those taking their driving test for the first time. What is the probability that among these 11 people, the number passing the test is between 2 and 4 inclusive?
Answer by stanbon(48568)

In one city, the probability that a person will pass his or her driving test on the first attempt is 0.67. 11 people are selected at random from among those taking their driving test for the first time. What is the probability that among these 11 people, the number passing the test is between 2 and 4 inclusive?
----------------
This is a binomial problem with n=11, p=0.67, 2<=X<=4
Using a TI-83 I get P(2<=X<=4)=0.36
You could also do this as P(X=2)+P(X=3)+P(X=4)
Cheers,
Stan H.